OpenStudy (anonymous):

I need help figuring this out: A certain elevator cab in a New York hotel has a total run of 160 m. Its maximum speed is 296 m/min. Its acceleration (both speeding up and slowing down) has a magnitude of 1.40 m/s2. How far does the cab move while accelerating to full speed from rest?

OpenStudy (anonymous):

T=(296-160)/1.4 X=V*T+.5*A*T^2 FINALLY X=22136.8